OpenStudy (anonymous):

solve. 10x^4-11x^2+3=0

OpenStudy (anonymous):

helps if put the whole equation in!

OpenStudy (saifoo.khan):

sorry, what do u mean?

OpenStudy (anonymous):

i wrote it without the"=0" a minute ago!

OpenStudy (saifoo.khan):

\[- \sqrt{\frac35}, \sqrt{\frac35}, - \frac{1}{\sqrt2}, \frac{1}{\sqrt2}\]

OpenStudy (saifoo.khan):

are u doing a final?

OpenStudy (anonymous):

trying to. wanting to double check my answers and some of them I am just confused on! have to get a good grade on this final to pass the class!

OpenStudy (saifoo.khan):

I hope u get an A. :D my highest score is 96. =D

OpenStudy (anonymous):

thank you! i need at least a C but think with all of the help I am getting plus what I actually understand, I might be able to do it!

OpenStudy (saifoo.khan):

wow, really.. if u want to any help, i can help you! ^_^

OpenStudy (anonymous):

I can use all of the help I can get! lol

OpenStudy (saifoo.khan):

lol, cool. =D